You plan to conduct a survey to find what proportion of the workforce has two or more jobs. You decide on the 98% confidence level and a margin of error of 2%. A pilot survey reveals that 7 of the 50 sampled hold two or more jobs. (Use z Distribution Table.)

How many in the workforce should be interviewed to meet your requirements?
Answer #1

p = 7/50 = 0.14, q = 1 - p = 1 - 0.14 = 0.86

For 98% confidence, z = 2.33

E = 0.02

Hence,

Sample size required

n = 1635